Given that 57% of all gold dealers believe next year will be a good one to speculate in south afrucan gold coins, in a simple random sample of 150 dealers, what is the probability that between 54% and 63% believe that it will be a good year to speculate? Please show work
Answer #1

Given. Sample size(n)= 150 Proportion (p=0.57 Implies. q=1-p = 1-0.57 = 0.43The required probability is. 0.54-0.570 - p P(0.54 < p <0.63) = P - 0.57) (0.43) V 150 Vn 0.63 -0.57 (0.57)(0.43) V 150 P 0.5